OpenStudy (anonymous):

Solve the Differential equation \[\frac{ dx }{ dy } = \frac{ x+y+7 }{ 2x+xy+3 }\]

OpenStudy (anonymous):

U/V = [dU(V) - dV(U)] / v^2 u = x + y + 7 v = 2x + xy +3 dU = derivative of x + y + 7 dV = derivative of 2x + xy +3

OpenStudy (perl):

he wants to integrate, not take derivative

OpenStudy (perl):

the derivative is given, what is the original function?

OpenStudy (unklerhaukus):

do you mean dx/dy ? or dy/dx?

OpenStudy (anonymous):

it is dx/dy

OpenStudy (anonymous):

oops. sorry

OpenStudy (perl):

this is a nonlinear differential equation, its tough. do you have any hints given

OpenStudy (anonymous):

Nothing..

OpenStudy (perl):

a substitution might work

OpenStudy (anonymous):

@satellite73 @phi

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@eliassaab @electrokid @Mashy @shubhamsrg

OpenStudy (anonymous):

@eliassaab @tkhunny

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

@satellite73 @helder_edwin

OpenStudy (anonymous):

@experimentX